Please help me solve this problem:
Five years ago, Dino was 1/5 as old as his father. Thirteen years from now, Dino will be half as old as his father. Find their present age.
Thanks..

Answer by mananth(16946)

Let dinos' age be x five years ago
father's age = 5x
After 13 years
dino's age = x+18
Father's age = 5x+18
2(x+18)= 5x+18
2x+36= 5x+18
5x-2x= 36-18
3x=18
/3
x= 6 years 5 years ago.
Now Dino's age = 5+5 = 10 years
Father's age 5 years ago = 5x= 30 years
Now Father's age = 35 years